5 cubic meter tank contains chlorine at 250 kPa and 30 degrees Celsius after 2.5 kg of chlorine has been used.

Determine: 

 

(a) the original mass of chlorine in the tank in kg

(b) the pressure (kPa) of chlorine in the tank if the original temperature was 315K


1
Expert's answer
2021-11-18T10:16:35-0500

Solution;

Given;

"V=5m^3"

"P=250kPa"

"T=30\u00b0c=303K"

Molecular weight of chlorine is 70.9

Assume that chlorine to an ideal gas;

"PV=mRT"

"R=\frac{\bar{R}}{M}=\frac{8.314}{70.9}"

"m=\frac{PV}{RT}=\frac{250×5}{\frac{8.314}{70.9}×303}" "m=\frac{1250}{35.53}=35.18kg"

(a)

Original mass;

"m_o=35.18+2.5=37.68kg"

(b)

"T=315K"

"PV=m_oRT"

"P=\frac{m_oRT}{V}=\frac{37.68×0.1173×315}{5}=278.37kPa"